OpenStudy (amatuluve):

Please help 1) f(x) = x^2 + 3; g(x) = x - 7 which expression is equal to (f + g)(x)? x^2-14x+52 x^2-4 x^3-7x^2+3x-21 x^2+x-4 2) f(x) = 4x - 1; g(x) = 5x^2 which expression is equal to (fog)(x)? 5x^2 + 4x - 1 20x^2 - 1 80x^2 - 40x + 5 20x^3-5x^2

OpenStudy (amorfide):

(f+g)(x) = f(x)+g(x)=(x^2 + 3) + (x - 7)

OpenStudy (amatuluve):

3) f(x) 3x-9; g(x) = x-4 what is the value of (gof)(5) 5 6 2 9 4) what is the value of f(-4)? f(x) = {-10 if x<-5 { x^3 if-5≤x≤2 { 2x+4 if x>2 -12 -64 -10 -4

OpenStudy (amorfide):

(fog)(x)=f((g(x)) which means let x=g(x) so substitute the entire g(x) function into f(x)

OpenStudy (amorfide):

(gof)(5)=g(f(5)) which means do f(5), where x= 5 then substitute this value into g(x) where x=f(5)

OpenStudy (amatuluve):

cant you just tell me what the answer is for all of them?

OpenStudy (amatuluve):

please help @zepdrix

OpenStudy (amorfide):

if you put some effort into attempting them then sure

OpenStudy (amatuluve):

I don't know how to d this?

OpenStudy (amorfide):

if I had f(x)=3x then said I want f(4) I would let x=4 to get f(4)=3(4) f(4)=12 so if I had f(x)=3x g(x)=5x then (f+g)(x)=f(x)+g(x)=3x+5x so this means I add the functions together if I have (fog)(x)=f(g(x)) we always start from right to left, so you do g(x) first, then substitute this into f(x) where x=g(x) we know that g(x)=5x so now we let x=5x I get (fog)(x)=f(g(x))=f(g(x)) f(5x) so f(5x)=3(5x)=15x if I had (gof)(x) we would do g(f(x)) instead

OpenStudy (mathmale):

The topic here is "composite functions." You might get more help, faster, if you would (1) post only one question at a time, (2) share whatever you have done on a given problem, and (3) not ask for answers without first having put some effort of your own into solving problems here.

OpenStudy (amatuluve):

Ok thanks!

OpenStudy (mathmale):

To obtain (f+g)(x), merely add f(x) and g(x) together. Try it.